When the smaller of two consectutive integers is added to five times the larger, the result is 47. Setting it up is fine I think I can solve it if it's set up right. TY

Answer by jorel1380(3719)

n+5(n+1)=47
n+5n+5=47
6n=42
n=7
n+1=8
The integers are 7 and 8..